OpenStudy (anonymous):

Forty-eight is what percent of 60? A. 80% B. 88% C. 120% D. 125%

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

answer pls

OpenStudy (anonymous):

im working on it for u dont worry

OpenStudy (anonymous):

No answers.

OpenStudy (anonymous):

IK

OpenStudy (anonymous):

first one to answer first get fanned and medalled

OpenStudy (anonymous):

and explained wins

OpenStudy (anonymous):

ITS A

OpenStudy (anonymous):

plan your out

OpenStudy (anonymous):

you have to explanied it

OpenStudy (nevermind_justschool):

Hey, this website is a place where you get help with your questions, not the answers. Learn more about the rules by reading: http://openstudy.com/code-of-conduct Please, please no direct answers.

OpenStudy (anonymous):

A- 80% Here is how you solve for it 48= n% x 60 n% = 48/60 x 100% n% = 80%

OpenStudy (anonymous):

Forty-eight is 80% of 60. Solution: 48/60 = x/100; 60x = 48(100); 60x = 4800; x = 4800/60; x = 80

OpenStudy (anonymous):

first u want to time 48 by 80

OpenStudy (anonymous):

and u will get 3840

OpenStudy (anonymous):

Forty-eight is 80% of 60. Solution: 48/60 = x/100; 60x = 48(100); 60x = 4800; x = 4800/60; x = 80

OpenStudy (anonymous):

then u divide 3840 by 80

OpenStudy (anonymous):

how did i geta metal?

OpenStudy (anonymous):

by explaining it like this Forty-eight is 80% of 60. Solution: 48/60 = x/100; 60x = 48(100); 60x = 4800; x = 4800/60; x = 80

OpenStudy (anonymous):

well thank u btw and ther ur answer with out me saying it directly

OpenStudy (anonymous):

im going to fan u

OpenStudy (anonymous):

ok the winner is

OpenStudy (anonymous):

?